# Navigating Cookie Consent Violations Across the Globe

### v1

@@ 1 Introduction
One of the most common ways of collecting data from users’ web browsers is through the use of cookies. Cookies are used to remember users’ preferences, transmit information such as location or IP addresses, and track users’ browsing history. However, the collection and use of personal data through cookies have raised significant privacy concerns. In response to these concerns, regions/countries worldwide have enacted laws requiring disclosure of privacy practices and/or user consent for data collection. For instance, the EU’s General Data Protection Regulation (GDPR) [gdpr_2016] and ePrivacy Directive (ePD) [epd_2002] suggest online services obtain consent before collecting personal data for advertising. Similarly, the California Consumer Privacy Act (CCPA) in the U.S., Canada’s Personal Information Protection and Electronic Documents Act (PIPEDA), and Singapore’s Personal Data Protection Act (PDPA), are among other similar privacy regulations that provide guidelines on consent, online behavioral advertising (OBA), and tracking cookies. These policies reflect a global shift towards enhanced data privacy protections.
To comply with these regulations, websites employ mechanisms and third-party services known as Consent Management Platforms (CMPs) to allow users to manage their cookie preferences. Figure. 1 illustrates an example of rejecting cookies from Google Analytics. Despite these measures, prior research indicates that cookies are often placed on browsers even after users have explicitly rejected them through cookie settings interfaces [bollinger2022automating, matte_cookie_2020]. These inconsistencies in cookie banner behavior circumvent users’ consent preferences and are known as cookie consent violations or cookie violations. Such practices not only undermine user trust but may also go against regional privacy guidelines [federal_trade_commission_goldenshores_2013, european_data_protection_board_belgian_2019, dataguidance_france_2021].
@@
ConsentChk addresses several technical challenges in detecting cookies that violate user consent. It detects and activates cookie banners’ menus, setting the consent choice for each cookie (accept/reject), and checks for inconsistencies in the expected and observed behavior. We conduct an evaluation analyzing cookie consent behavior for the same set of websites across different regions with different regulatory privacy frameworks. Using ConsentChk, we have examined 1,793 websites from the 20k top global websites according to the Tranco list [LePochat_tranco_2019].
@@
  • An end-to-end measurement tool, called ConsentChk, that crawls websites, detecting infringements on users’ consent preferences for each cookie on a website. ConsentChk uses an approach that identifies features of cookie banner buttons to activate the cookie banner menus of any website automatically. The detection of cookie consent violations has a high precision of >91%. (Sections 4 and 4.6)
  • A study on 1,793 of the top visited websites in the globe based on the Tranco 1M list [LePochat_tranco_2019] with cookie banners. ConsentChk detected that 96.18% (EU) – 97.72% (US) of websites across all regions contain at least one cookie consent violation. These violations primarily occur in the form of undeclared cookies or ignored cookie rejections, demonstrating that cookie banners and CMPs have inadequate coverage and correctness. (Sections 5 and 6)
  • The discovery of discrepancies in cookie consent behavior across various regions. We observe that, compared to the EU, the same websites in the US have an average of almost 12 additional cookies, and an average of 10 additional cookie consent violations. Similarly, regions like Canada, Singapore, South Africa, and Australia with their own privacy frameworks have 6-8 more cookies and 5-7 more cookie consent violations than the EU. (Section 6)
  • An analysis of the potential root causes behind cookie consent violations and regional discrepancies. CMPs provide banner customization and geolocation rulesets. Thus, cookie banner buttons and cookie behaviors are largely reliant on factors such as location, advertiser interest, cookie scraping/classification accuracy, and more. (Section 7)
@@ 2 Background
CAN PIPEDA Schedule 1, Principle 3 (4.3.8), Sec 1.6.1 Yes [canada2017courtruling] Informed Info About Person, Includes Cookies, Identifiers
@@ 2.1 Privacy Laws Across the Globe
GDPR (European Union, EU). Consent is one of 6 lawful bases defined in the GDPR for processing data. The others include a contractual obligation, a legal obligation, vital interests, a task in public interest, and legitimate interests. Many websites provide a privacy policy or notice for the data they process under these other 5 bases (for purposes such as marketing, fraud prevention, etc.), but additionally provide a cookie banner for obtaining consent for the usage of cookies. Article 7 of the General Data Protection Regulation (GDPR) discusses informed consent related to the processing of users’ personal data. In particular, Article 7.3 states that data subjects can withdraw their consent at any time and that it should be as easy to withdraw as to give consent [eu:gdpr].
@@
DPA (United Kingdom, UK). The UK’s Data Protection Act (DPA) is similar to GDPR, following and adapting the same protection rules to the UK’s legal system [cookiebot_new_2021, information_commissioners_office_uk_2022, part2018dpa].
@@
CCPA (California, CA). Some privacy laws, such as the California Consumer Privacy Act (CCPA) do not require consent for data processing. Unlike the GDPR, a lawful basis for processing data is not required, however, businesses cannot process data in unfair or deceptive manners. Instead, this framework requires that businesses meeting certain criteria (large amount of revenue from selling personal information, with certain exemptions) give consumers information about data collection activities and purposes in a “notice at collection” [ccpa2023], giving users the right to access, delete, and opt out of the selling of their data.
@@
PIPEDA (Canada, CAN). Canada’s Personal Information Protection and Electronic Documents Act (PIPEDA’s) applies to organizations that collect, use, or disclose personal information in the course of a commercial activity. Fair Information Principle 3 states organizations must generally obtain express consent when the information being collected, used, or disclosed is sensitive [pipeda2025, pipeda2021tracking].
@@
POPIA (South Africa, ZA). Similarly to the GDPR, the Protection of Personal Information Act (POPIA) identifies 5 legal bases for processing data, including consent from the data subject. It requires organizations to obtain consent before collecting and processing personal information. Additionally, data subjects have the right to be notified of data collection, request correction, destruction, or deletion, and to object to processing of data for marketing [popia2019, popia2019consent].
@@
PDPA (Singapore, SG). The Personal Data Protection Act (PDPA) states in Part 4, Division 1 that an organization must not use or disclose personal data about an individual unless the individual gives, or is deemed to have given, his/her consent to the collection, use, or disclosure. It also states that users may at any time withdraw any consent given in respect of the collection, use, or disclosure by that organization of personal data about the individual [pdpa2012].
@@
APP (Australia, AU). Section 5.1 of the Australian Privacy Principles (APP) requires entities to disclose purposes and request consent at or before the time or, if that is not practicable, as soon as practicable after, collecting personal information about an individual. Section 6 requires entities to not use or disclose the collected personal information for another purpose without consent [app2025, app2024cookies].
@@ 2.2 Role of Cookie Consent in Privacy Laws
While consent is one of many frameworks for allowing users to control and understand privacy practices, it is the primary form of control used for cookies on websites, due to the prevalence of cookies in online behavioral advertising. In Europe, the ePrivacy Directive (ePD) was an EU directive which modernized privacy laws to apply to cookies [epd_2002]. Cookies (other than those used for strictly necessary services) required user consent. These new guidelines paved the way for the IAB’s Transparency and Consent Framework (TCF) to be created [iab_europe_tcf_2021]. The TCF is a voluntary standard providing publishers (websites) and advertisers (vendors) with tools and libraries for providing and managing consent options for users.
@@
Due to the extra-territoriality clause of privacy laws like GDPR and CCPA, companies are required to comply with laws so long as they do business with any EU or CA resident. Website developers cannot be expected to handle privacy compliance with so many regions outside of their own company’s location. As such, intermediary 3rd-party consent management platforms (CMPs) have standardized compliance for website developers. Based on the provisions put forth in the various regions worldwide, these CMPs convert interpretations of privacy laws into implementations of privacy laws. These CMPs have standardised frameworks for regional privacy laws, directly impacting website developers’ cookie banner implementations. CMPs manage users’ consent, which are important for compliance with GDPR in Europe, PIPEDA in Canada, CCPA in California, and more. Hence, CMPs are often implemented such that the geolocation from a user’s browsing session informs the CMP to display a particular cookie banner with its UI and behavior customized to the region’s regulations [onetrust2025regulations]. For example, in Europe, users more frequently encounter detailed consent banners with guidelines set by the GDPR, ePD, and TCF, whereas in the U.S., banners may only notify users of data collection practices, or simply not appear at all.
@@ 2.3 Related Work
Browser extensions such as Consent-O-Matic and Autoconsent automate rejection by matching CSS/JS patterns. These programs use handcrafted CSS matchers and JavaScript object models to create a JavaScript model of consent notices and automatically set the cookie banners to rejection. Ad blockers’ “annoyance blocking” closes/hides banners but does not actively reject cookies. Sweepatic’s platform detects non-consensual cookies but only before a consent choice being made [sweepatic_sweepatic_2022]. Several researchers study the effectiveness of these automated opt-outs such as Khandelwal et al. [khandelwal_cookieenforcer_2022] which model consent dialogs as seq2seq tasks and Demir et al. [demir2024large] which compares 6 tools over 98k pages from 30k sites.
@@
Cookie compliance has been studied since the GDPR’s enforcement in 2018. Sanchez-Rola et al. [sanchez-rola_can_2019] manually analyzed 2,000 sites across EU, USA, China, and other regions, reporting that 92% perform tracking before consent and only 2.5% erase cookies after their placement. They also report that cookie-settings interfaces appear on just 16% of EU sites and 12% of US sites. Matte et al. [matte_cookie_2020] also analyzed 1,426 EU sites using IAB TCF banners, finding that 54% contained violations like pre-selected options, consent stored pre-choice, etc. These findings were extended to 3rd-party cookies as well by Kancherla et al. [kancherla2024johnny], who inspected 1,200 sites (top 200 for UI, 1,000 random from top 5k), reporting that 74% do not inform third parties of rejection. Cookies that remain on the browser were also found to be continually used even after rejection. Liu et al. [liu_opted_2022] compared advertiser bidding before and after GDPR/CCPA opt-out (across 4 CMPs and 34 IPs) and found negligible differences in tracking.
@@
Bollinger et al. [bollinger2022automating] and Bouhoula et al. [bouhoula2023automated] focused on CMP compliance, defining 8 violation types (e.g., undeclared cookies and implicit consent) and scaled up compliance testing to 30k and 97k sites in the EU, respectively. The former found at least one violation in 94.7% of sites, with 82.5% sites containing undeclared cookies and 69.7% sites with implicit consent, and the latter similarly found 65.4% were likely to collect data despite negative consent.
@@
Measurement studies have also been conducted on cookie banners in regions world-wide. Van Eijk et al. [eijk_impact_2019] crawled 1,543 sites from 18 countries (27,488 measurements in total) and found that the odds of seeing a cookie banner are 102% higher in the EU. Rasaii et al. [rasaii2023exploring] examined 518 sites across 8 locations (Sweden, Germany, US West/East, Brazil, South Africa, Australia, India), finding cookie banners in 47% in EU vs. 30% non-EU.
@@
Santos et al. [santos2019cookie] identified 22 legal requirements for valid consent (e.g., free, specific, informed, prior, revocable) and showed that no CMP satisfies them all. Santos et al. [santos2021consent] also determined that CMPs act as both data controllers/processors, noting that HTTPS-based CMP s cripts process IP addresses (an identifier) and shape users’ choices. Bouma-Sims et al. [bouma2023us] performed a usability study of OneTrust’s UI across US and UK participants, finding that over half could not make a properly informed consent decision (52.4% US vs. 46.2% UK).
@@
Unlike prior studies that focus on a single region, law, or CMP, we measure cookie consent behavior across 8 jurisdictions, 3 major CMPs, and websites from the Tranco top 20k. By analyzing legal regulations and CMP documentation, we identified the root causes that result in cookie banner/consent/placement disparities and violations. We conducted measurements at the individual cookie level, identifying cookies that likely contained personal information. ConsentChk is the first to compare cookie consent violations across multiple non-EU regions with privacy laws. Table 2 provides a summary.
@@ 3 Definitions and Terminology
We provide formal definitions of cookies, consent preferences, and cookie consent violations in Table 3.
@@
“Cookie banners” inform users of the use of cookies (providing no option or confirmation only) and allow users to accept, reject, or manage their cookie consent preferences [degeling_we_2019]. Cookie banners can appear as pop-ups on landing pages or can be activated by buttons placed in footers or cookie/privacy policy pages. These banners are often hidden and must be manually activated in regions where privacy laws do not require consent before data collection. Oftentimes, these banners allow users to open separate “cookie preference menus” with controls that allow users to set their consent preference on specific cookie categories or individual cookies (Figure. 1). The categories can be specified by purposes and/or by advertising vendors.
Websites collect users’ cookie consent preferences and block/unblock cookies locally based on the consent choice. CMPs create and maintain special cookies to record users’ consent choices on websites. For example, OneTrust and Cookiebot store users’ consent preference in cookies named OptanonConsent [onetrust_llc_onetrust_2021] and CookieConsent [cybot_duration_2020], respectively. Other sites use 1st-party cookie banners instead of CMPs.
@@
Considering a cookie in a browser’s cookie storage as a tuple of key–values, ConsentChk distinguishes cookies by their name, domain, and path. This distinction of cookies follows the storage model in the cookie specification [barth_http_2011]. Other cookie attributes, while still important to monitor, may change over time even for the same cookie (e.g., value and expiration time).
@@
We selected three main forms of cookie consent violations to investigate in this paper. These violations are defined as when cookie consent mechanisms are not consistent in their expected/specified behavior and may mislead users to believing that tracking cookies are not placed on their browsers. These three violations are originally derived by us but are similar in nature and scope to the cookie consent violations explored in Bollinger et al.’s work [bollinger2022automating].
@@
A Wrong Cookie Category violation is when the cookie’s consent is accepted and rejected at the same time. This occurs when the categories of cookie banners overlap, allowing a cookie to be both rejected and approved by a user. For example, a cookie could be included in both the Strictly Necessary and Targeting categories. A user could have rejected targeting cookies but accepted strictly necessary cookies. This is akin to contradictory statements in privacy policies identified by the FTC and prior studies [PolicyLint_2019, purpliance_bui_2021, pitofsky_privacy_2000].
The existence of Ignored Cookie Rejections and Undeclared Cookies seemingly violate the GDPR, as Recital 32 states that the conditions for consent require that “consent should not be regarded as freely given if the data subject has no genuine or free choice or is unable to refuse or withdraw consent without detriment”. In Canada, PIPEDA guidelines state that consent is reasonable for online behavioral advertising if individuals are aware, informed of these practices and are easily able to opt out. The guidelines also state “If an individual is not able to decline the tracking and targeting…then organizations should not be employing that type of technology for online behavioral advertising [pipeda2021tracking].” Similar articles exist in other privacy frameworks like PDPA and others.

### v2

@@ 1 Introduction
One of the most common ways of collecting data from users’ web browsers is through the use of cookies. Cookies are used to remember users’ preferences, transmit information such as location or IP addresses, and track users’ browsing history. However, the collection and use of personal data through cookies have raised significant privacy concerns. In response to these concerns, regions/countries worldwide have enacted laws requiring disclosure of privacy practices and/or user consent for data collection. For instance, the EU’s General Data Protection Regulation (GDPR) [65] and ePrivacy Directive (ePD) [64] suggest online services obtain consent before collecting personal data for advertising. Similarly, the California Consumer Privacy Act (CCPA) in the U.S., Canada’s Personal Information Protection and Electronic Documents Act (PIPEDA), and Singapore’s Personal Data Protection Act (PDPA), are among other similar privacy regulations that provide guidelines on consent, online behavioral advertising (OBA), and tracking cookies. These policies reflect a global shift towards enhanced data privacy protections.
To comply with these regulations, websites employ mechanisms and third-party services known as Consent Management Platforms (CMPs) to allow users to manage their cookie preferences. Figure. 1 illustrates an example of rejecting cookies from Google Analytics. Despite these measures, prior research indicates that cookies are often placed on browsers even after users have explicitly rejected them through cookie settings interfaces [9, 45]. These inconsistencies in cookie banner behavior circumvent users’ consent preferences and are known as cookie consent violations or cookie violations. Such practices not only undermine user trust but may also go against regional privacy guidelines [16, 8, 25].
@@
ConsentChk addresses several technical challenges in detecting cookies that violate user consent. It detects and activates cookie banners’ menus, setting the consent choice for each cookie (accept/reject), and checks for inconsistencies in the expected and observed behavior. We conduct an evaluation analyzing cookie consent behavior for the same set of websites across different regions with different regulatory privacy frameworks. Using ConsentChk, we have examined 1,793 websites from the 20k top global websites according to the Tranco list [42].
@@
  • An end-to-end measurement tool, called ConsentChk, that crawls websites, detecting infringements on users’ consent preferences for each cookie on a website. ConsentChk uses an approach that identifies features of cookie banner buttons to activate the cookie banner menus of any website automatically. The detection of cookie consent violations has a high precision of >91%. (Sections˜4 and 4.6)
  • A study on 1,793 of the top visited websites in the globe based on the Tranco 1M list [42] with cookie banners. ConsentChk detected that 96.18% (EU) – 97.72% (US) of websites across all regions contain at least one cookie consent violation. These violations primarily occur in the form of undeclared cookies or ignored cookie rejections, demonstrating that cookie banners and CMPs have inadequate coverage and correctness. (Sections˜5 and 6)
  • The discovery of discrepancies in cookie consent behavior across various regions. We observe that, compared to the EU, the same websites in the US have an average of almost 12 additional cookies, and an average of 10 additional cookie consent violations. Similarly, regions like Canada, Singapore, South Africa, and Australia with their own privacy frameworks have 6-8 more cookies and 5-7 more cookie consent violations than the EU. (Section˜6)
  • An analysis of the potential root causes behind cookie consent violations and regional discrepancies. CMPs provide banner customization and geolocation rulesets. Thus, cookie banner buttons and cookie behaviors are largely reliant on factors such as location, advertiser interest, cookie scraping/classification accuracy, and more. (Section˜7)
@@ 2 Background
CAN PIPEDA Schedule 1, Principle 3 (4.3.8), Sec 1.6.1 Yes [1] Informed Info About Person, Includes Cookies, Identifiers
@@ 2.1 Privacy Laws Across the Globe
GDPR (European Union, EU). Consent is one of 6 lawful bases defined in the GDPR for processing data. The others include a contractual obligation, a legal obligation, vital interests, a task in public interest, and legitimate interests. Many websites provide a privacy policy or notice for the data they process under these other 5 bases (for purposes such as marketing, fraud prevention, etc.), but additionally provide a cookie banner for obtaining consent for the usage of cookies. Article 7 of the General Data Protection Regulation (GDPR) discusses informed consent related to the processing of users’ personal data. In particular, Article 7.3 states that data subjects can withdraw their consent at any time and that it should be as easy to withdraw as to give consent [30].
@@
DPA (United Kingdom, UK). The UK’s Data Protection Act (DPA) is similar to GDPR, following and adapting the same protection rules to the UK’s legal system [18, 48, 50].
@@
CCPA (California, CA). Some privacy laws, such as the California Consumer Privacy Act (CCPA) do not require consent for data processing. Unlike the GDPR, a lawful basis for processing data is not required, however, businesses cannot process data in unfair or deceptive manners. Instead, this framework requires that businesses meeting certain criteria (large amount of revenue from selling personal information, with certain exemptions) give consumers information about data collection activities and purposes in a “notice at collection” [14], giving users the right to access, delete, and opt out of the selling of their data.
@@
PIPEDA (Canada, CAN). Canada’s Personal Information Protection and Electronic Documents Act (PIPEDA’s) applies to organizations that collect, use, or disclose personal information in the course of a commercial activity. Fair Information Principle 3 states organizations must generally obtain express consent when the information being collected, used, or disclosed is sensitive [52, 33].
@@
POPIA (South Africa, ZA). Similarly to the GDPR, the Protection of Personal Information Act (POPIA) identifies 5 legal bases for processing data, including consent from the data subject. It requires organizations to obtain consent before collecting and processing personal information. Additionally, data subjects have the right to be notified of data collection, request correction, destruction, or deletion, and to object to processing of data for marketing [41, 40].
@@
PDPA (Singapore, SG). The Personal Data Protection Act (PDPA) states in Part 4, Division 1 that an organization must not use or disclose personal data about an individual unless the individual gives, or is deemed to have given, his/her consent to the collection, use, or disclosure. It also states that users may at any time withdraw any consent given in respect of the collection, use, or disclosure by that organization of personal data about the individual [51].
@@
APP (Australia, AU). Section 5.1 of the Australian Privacy Principles (APP) requires entities to disclose purposes and request consent at or before the time or, if that is not practicable, as soon as practicable after, collecting personal information about an individual. Section 6 requires entities to not use or disclose the collected personal information for another purpose without consent [55, 63].
@@ 2.2 Role of Cookie Consent in Privacy Laws
While consent is one of many frameworks for allowing users to control and understand privacy practices, it is the primary form of control used for cookies on websites, due to the prevalence of cookies in online behavioral advertising. In Europe, the ePrivacy Directive (ePD) was an EU directive which modernized privacy laws to apply to cookies [64]. Cookies (other than those used for strictly necessary services) required user consent. These new guidelines paved the way for the IAB’s Transparency and Consent Framework (TCF) to be created [29]. The TCF is a voluntary standard providing publishers (websites) and advertisers (vendors) with tools and libraries for providing and managing consent options for users.
@@
Due to the extra-territoriality clause of privacy laws like GDPR and CCPA, companies are required to comply with laws so long as they do business with any EU or CA resident. Website developers cannot be expected to handle privacy compliance with so many regions outside of their own company’s location. As such, intermediary 3rd-party consent management platforms (CMPs) have standardized compliance for website developers. Based on the provisions put forth in the various regions worldwide, these CMPs convert interpretations of privacy laws into implementations of privacy laws. These CMPs have standardised frameworks for regional privacy laws, directly impacting website developers’ cookie banner implementations. CMPs manage users’ consent, which are important for compliance with GDPR in Europe, PIPEDA in Canada, CCPA in California, and more. Hence, CMPs are often implemented such that the geolocation from a user’s browsing session informs the CMP to display a particular cookie banner with its UI and behavior customized to the region’s regulations [60]. For example, in Europe, users more frequently encounter detailed consent banners with guidelines set by the GDPR, ePD, and TCF, whereas in the U.S., banners may only notify users of data collection practices, or simply not appear at all.
@@ 2.3 Related Work
Browser extensions such as Consent-O-Matic and Autoconsent automate rejection by matching CSS/JS patterns. These programs use handcrafted CSS matchers and JavaScript object models to create a JavaScript model of consent notices and automatically set the cookie banners to rejection. Ad blockers’ “annoyance blocking” closes/hides banners but does not actively reject cookies. Sweepatic’s platform detects non-consensual cookies but only before a consent choice being made [62]. Several researchers study the effectiveness of these automated opt-outs such as Khandelwal et al. [39] which model consent dialogs as seq2seq tasks and Demir et al. [27] which compares 6 tools over 98k pages from 30k sites.
@@
Cookie compliance has been studied since the GDPR’s enforcement in 2018. Sanchez-Rola et al. [56] manually analyzed 2,000 sites across EU, USA, China, and other regions, reporting that 92% perform tracking before consent and only 2.5% erase cookies after their placement. They also report that cookie-settings interfaces appear on just 16% of EU sites and 12% of US sites. Matte et al. [45] also analyzed 1,426 EU sites using IAB TCF banners, finding that 54% contained violations like pre-selected options, consent stored pre-choice, etc. These findings were extended to 3rd-party cookies as well by Kancherla et al. [38], who inspected 1,200 sites (top 200 for UI, 1,000 random from top 5k), reporting that 74% do not inform third parties of rejection. Cookies that remain on the browser were also found to be continually used even after rejection. Liu et al. [43] compared advertiser bidding before and after GDPR/CCPA opt-out (across 4 CMPs and 34 IPs) and found negligible differences in tracking.
@@
Bollinger et al. [9] and Bouhoula et al. [10] focused on CMP compliance, defining 8 violation types (e.g., undeclared cookies and implicit consent) and scaled up compliance testing to 30k and 97k sites in the EU, respectively. The former found at least one violation in 94.7% of sites, with 82.5% sites containing undeclared cookies and 69.7% sites with implicit consent, and the latter similarly found 65.4% were likely to collect data despite negative consent.
@@
Measurement studies have also been conducted on cookie banners in regions world-wide. Van Eijk et al. [28] crawled 1,543 sites from 18 countries (27,488 measurements in total) and found that the odds of seeing a cookie banner are 102% higher in the EU. Rasaii et al. [54] examined 518 sites across 8 locations (Sweden, Germany, US West/East, Brazil, South Africa, Australia, India), finding cookie banners in 47% in EU vs. 30% non-EU.
@@
Santos et al. [57] identified 22 legal requirements for valid consent (e.g., free, specific, informed, prior, revocable) and showed that no CMP satisfies them all. Santos et al. [58] also determined that CMPs act as both data controllers/processors, noting that HTTPS-based CMP s cripts process IP addresses (an identifier) and shape users’ choices. Bouma-Sims et al. [11] performed a usability study of OneTrust’s UI across US and UK participants, finding that over half could not make a properly informed consent decision (52.4% US vs. 46.2% UK).
@@
Unlike prior studies that focus on a single region, law, or CMP, we measure cookie consent behavior across 8 jurisdictions, 3 major CMPs, and websites from the Tranco top 20k. By analyzing legal regulations and CMP documentation, we identified the root causes that result in cookie banner/consent/placement disparities and violations. We conducted measurements at the individual cookie level, identifying cookies that likely contained personal information. ConsentChk is the first to compare cookie consent violations across multiple non-EU regions with privacy laws. Table˜2 provides a summary.
@@ 3 Definitions and Terminology
We provide formal definitions of cookies, consent preferences, and cookie consent violations in Table˜3.
@@
“Cookie banners” inform users of the use of cookies (providing no option or confirmation only) and allow users to accept, reject, or manage their cookie consent preferences [26]. Cookie banners can appear as pop-ups on landing pages or can be activated by buttons placed in footers or cookie/privacy policy pages. These banners are often hidden and must be manually activated in regions where privacy laws do not require consent before data collection. Oftentimes, these banners allow users to open separate “cookie preference menus” with controls that allow users to set their consent preference on specific cookie categories or individual cookies (Figure. 1). The categories can be specified by purposes and/or by advertising vendors.
Websites collect users’ cookie consent preferences and block/unblock cookies locally based on the consent choice. CMPs create and maintain special cookies to record users’ consent choices on websites. For example, OneTrust and Cookiebot store users’ consent preference in cookies named OptanonConsent [49] and CookieConsent [24], respectively. Other sites use 1st-party cookie banners instead of CMPs.
@@
Considering a cookie in a browser’s cookie storage as a tuple of key–values, ConsentChk distinguishes cookies by their name, domain, and path. This distinction of cookies follows the storage model in the cookie specification [6]. Other cookie attributes, while still important to monitor, may change over time even for the same cookie (e.g., value and expiration time).
@@
We selected three main forms of cookie consent violations to investigate in this paper. These violations are defined as when cookie consent mechanisms are not consistent in their expected/specified behavior and may mislead users to believing that tracking cookies are not placed on their browsers. These three violations are originally derived by us but are similar in nature and scope to the cookie consent violations explored in Bollinger et al.’s work [9].
@@
A Wrong Cookie Category violation is when the cookie’s consent is accepted and rejected at the same time. This occurs when the categories of cookie banners overlap, allowing a cookie to be both rejected and approved by a user. For example, a cookie could be included in both the Strictly Necessary and Targeting categories. A user could have rejected targeting cookies but accepted strictly necessary cookies. This is akin to contradictory statements in privacy policies identified by the FTC and prior studies [4, 12, 53].
The existence of Ignored Cookie Rejections and Undeclared Cookies seemingly violate the GDPR, as Recital 32 states that the conditions for consent require that “consent should not be regarded as freely given if the data subject has no genuine or free choice or is unable to refuse or withdraw consent without detriment”. In Canada, PIPEDA guidelines state that consent is reasonable for online behavioral advertising if individuals are aware, informed of these practices and are easily able to opt out. The guidelines also state “If an individual is not able to decline the tracking and targeting…then organizations should not be employing that type of technology for online behavioral advertising [33].” Similar articles exist in other privacy frameworks like PDPA and others.